OpenStudy (anonymous):

An advertising firm has an old computer that can prepare a whole mailing in 6 hours. With the help of a newer model the job is complete in 2 hours. How long would it take the newer model to do the job alone?

OpenStudy (anonymous):

I got 1/6 +1/x = 1/2 then 1/x = 1/2 -1/6 = 1/3

OpenStudy (anonymous):

I got 1/x = 1/3 what's next?

ganeshie8 (ganeshie8):

your goal is to solve \(x\)

OpenStudy (anonymous):

Yeap I know but how would I translate 1/x to x=3 I know 3 is the answer but how will I get there

ganeshie8 (ganeshie8):

1/x = 1/3 flip both sides and get x/1 = 3/1 which is same as x = 3

OpenStudy (anonymous):

You just flip it? No math explanation?

ganeshie8 (ganeshie8):

flipping is mathematical. for nonzero real numbers, we have : \[\dfrac{a}{b} = \dfrac{c}{d} \implies \dfrac{b}{a}=\dfrac{d}{c}\]

OpenStudy (anonymous):

Oh I thought when you flip it becomes \[x^-1 = 3\]

OpenStudy (anonymous):

wait a minute I could've use cross multiply!!!

ganeshie8 (ganeshie8):

Yes you could do that too

OpenStudy (anonymous):

Yeah thank you for lending me a hand. I'm doing crash course for my engineering math exit. Thank you!

ganeshie8 (ganeshie8):

np, good luck with the course!